A hot tungsten wire filament has a radius of 0.6cm a temperature of 300K and emissivity

0.74. Calculate the rate of energy emission by a 1.0m length of the wire. Ignore the radiation

received from the surrounding environment.


1
Expert's answer
2021-10-07T08:48:25-0400

The radiation heat current ("H") from a body at temperature T (T = 300 K) to its surroundings depends on the mentioned property, the surface area A ("A=2\pi r^2+2\pi rL=2\pi r(r+L)", that formula is the whole area of the exterior of the wire filament that has a radius r = 0.06 m and length L = 1 m), the emissivity e (e=0.74) and the Stefan-Boltzmann constant ("\sigma = 5.6704\times10^{-8}\frac{W}{m^2 \cdot K^4}" ).


We proceed to find "H" after substitution:


"H=Ae\sigma T^4=2\pi r(r+L) e\sigma T^4
\\ H=2\pi [(0.06\,m)(1+0.06)m](0.74)(5.6704\times10^{-8}\frac{W}{m^2 \cdot K^4})(300\,K)^4
\\ H=135.821\,W"


In conclusion, we find that the rate of loss of heat of the hot tungsten wire filament is about 135.821 W.

Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.